i have a question that involves a right triangle with only the hypostenuse number given. c=14
the question ask me to find the area and the value of x but they only gave me one number.. can you please help me

Answer by ewatrrr(24785)


Hi
Please repost with more detail as to the 'type' of right triangle,
for example, is the question involving an isosoles right triangle,
where the legs are of equal length?
for ex, if that were the case:
x^2 + x^2 = 14^2
2x^2 = 196
x^2 = 98
area of this rightΔ = (1/2) x*x = (1/2)x^2 = (1/2)98 = 49